# RACP: Risk-Aware Contingency Planning with Multi-Modal Predictions

### v1

@@ II-A1 Planning with multi-modal predictions
Another line of work for planning under uncertainty is by formulating the planning problem as a partially observable Markov decision process (POMDP) by constructing a belief tree based on a discrete set of obstacle vehicle’s intentions [28, 29]. Solving such problems, however, becomes computationally intractable when the problem size scales. To address this problem, multi-policy decision making [33, 34, 35] decomposes the belief tree into a limited number of closed-loop policies by leveraging semantic information. However, these approaches solve for the best ego-policy over all possible future realizations resulting in overly-conservative plans that do not exploit the multi-modality in the obstacle-vehicle behavior.
This issue can be tackled by separating the planning problem into short-term and long-term responses [14, 15, 17, 16]. This is realized by generating a short-term trajectory, for $t<t_{b}$, where $t_{b}$ indicates a branching time, which guarantees safety with respect to all possible future intents that the other agents can exhibit. After $t_{b}$ duration, it is assumed that the ego vehicle will be able to determine the intent of the human driver, and thus it is sufficient to safeguard against the most likely intent. Thus, the contingent plan can ensure safety only when the ego vehicle acquires a clear understanding of the intentions of other agents by the time of branching. However, in these works, the probability associated with the uncertain intents of the dynamic agents is assumed to be fixed, and the planning is executed in an open-loop fashion. Thus, an exact estimation of the branching time $t_{b}$ is required, otherwise a collision may occur [17]. Learning-based approaches have been recently introduced in the contingency planning context [36, 37], however, they are not interpretable and hard to tune.
@@ V-A3 Probabilistic Risk Assessment
Given the probability density function of the prediction model as indicated in (2), the probability of collision can be calculated by estimating the spatio-temporal overlap of the predicted modes with the ego-vehicle’s plan $\tau$,
@@
which is an integral of the Mixture of Gaussians (MoG) probability density function over a specified domain where the integration domain $D$ is defined as a circle whose center is located at the predicted vehicle pose $\boldsymbol{x}_{k}^{d}$ at stage $k$ along the prediction horizon, and its radius $r$ is the sum of the vehicle and obstacle radii. Here, we approximate the vehicle by two discs and calculate the probability of collision for each disc.
@@
After calculating the risk for each obstacle along the ego-vehicle’s planned trajectory, the predicted risk $\eta$ at the current time step is defined by maximizing risk for all obstacles, $o\in\mathcal{I}_{o}$, at all planning horizon stages,
@@
Figure 3: Snapshots from the CommonRoad simulation framework for an overtaking scenario where the ego-vehicle is depicted by the car icon and the obstacle-vehicle is illustrated as a blue rectangle. The obstacle-vehicle has three policies that it can execute, Λ={maintain speed,slow down,lane $\Lambda=\{\texttt{maintain speed},\texttt{slow down},\texttt{lane change}\}$, where the obstacle-vehicle’s predicted trajectories are illustrated in distinct colors in 2(a). We first sample a set of short-term plans, $\mathcal{T}_{0:t_{b}}$ visualized in black, till the branching time $t_{b}$ shown in 2(a). We then sample a set of long-term plans, $\mathcal{T}_{t_{b}:T}$ depicted in purple, conditioned on the terminal states of the short-term plans, shown in 2(b). For the long-term plans, a cost is assigned to each plan per obstacle-vehicle policy $\lambda_{i}$. In Fig. 2(c), the total cost of the entire plan is computed by (13a), and the total optimal plan is the one given by $\tau^{*}=\tau^{*}_{0:t_{b}}\cup\{\tau^{*}_{t_{b}:T}(\lambda_{1}),\tau^{*}_{t_{b}:T}(\lambda_{2}),\tau^{*}_{t_{b}:T}(\lambda_{3})\}$. In this example, the belief over the first policy is dominant, $b(\lambda_{1})=0.63$, compared to the rest causing the optimal short-term plan $\tau^{*}_{0:t_{b}}$ to be more biased towards $\tau^{*}_{t_{b:T}}(\lambda_{1})$.
@@ V-B Planner Design
Here it is crucial to emphasize that an advantage of using a sampling-based planner in the Frenet frame as a basis for our contingency planning framework is that the sampled trajectories cover various maneuvers that the ego-vehicle can have. Thus, in contrast to the approaches proposed in [30, 31, 32], our approach eliminates the necessity for a pre-established trajectory tree or branching topology to formulate contingent plans. Subsequently, the ego-vehicle is not confined to a specific set of predefined policies. This concept is further explained in Fig. 3 where we show how the optimal plan is selected from the sampled short-term and long-term trajectories. Additionally, it illustrates how the short-term plan is shaped based on the belief the ego-vehicle maintains over the obstacle-vehicle policies.
Fig. 5 shows how the belief the ego-vehicle maintains over the long-term plans affects the behavior of the short-term plan. It is observed that when the ego-vehicle has a higher belief in one of the human’s intents, the shared plan tends to be biased towards the corresponding contingent plan. In particular, as illustrated in Fig. 5-left, when the ego-agent has a higher belief that the human-driven vehicle aims at executing the lane-keeping policy, the shared-plan biases its motion to accelerate along its lane. On the other hand, when the higher belief is assigned to the lane-change maneuver, the shared-plan tends towards steering a bit to the right while decelerating. This is achieved by balancing the cost of the shared-plan according to the likelihood of the beliefs.
@@ VI-A Experimental Setup
In the second scenario, we show how our approach, unlike [30], can be generalized to multi-vehicle scenarios. This is illustrated in a T-junction scenario, where the ego-vehicle approaches an intersection together with two fellow traffic participants whose intents are not known to the ego-agent a prior  [^1] A video of the simulated experiments accompanies this paper.. The computer running the simulations is equipped with an Intel® CoreTM i7 CPU@2.6GHz.
@@ VI-B1 Obstacle-vehicle Model
Non-cont. MLE [48] 3.15 (0.396) 17.83 (0.51) 0.48 (0.12) 2.74 % 0.1081 5.278 (6.13)
Contingent w/o belief updater [15] 3.79 (0.281) 16.91 (0.46) 0.84 (0.08) 0.00 % 0.0462 2.693 (4.63)
@@ VI-B3 Qualitative Results
Fig. 7 shows the evolution of the ego-vehicle’s belief over the obstacle-vehicle policies over time. At the beginning, the ego-vehicle reveals its lane-change intention by swerving into the obstacle’s vehicle lane. Due to the reactive behavior of the obstacle-vehicle, the probability of it executing a lane-change in the ego-vehicle’s lane drops. However, the ego-vehicle could not complete the lane change since no valid trajectory is obtained that does not violate the safety constraints, and thus returns back to its original lane. Since the lane-change policy of the obstacle-vehicle, $b_{3}$, becomes relatively low, the ego-vehicle then initiates another attempt to overtake the obstacle vehicle which probes it to decelerate allowing the ego-vehicle to complete the lane change.
On the other hand, it was observed that the branch-MPC [30] brakes strangely while executing the lane-change maneuver relying on the obstacle-vehicle to yield to the ego-vehicle. However, such a maneuver is risky since it may result in collisions if the obstacle-vehicle does not react on time.
Figure 7: Belief evolution over obstacle vehicle’s policies along the timesteps. $b_{1},b_{2},b_{3}$ represent maintain fixed speed, slow down, and lane-change policies respectively.
@@ VI-B4 Quantitative Results
Figure 8: Snapshots from the T-junction simulated environment in CommonRoad at different time instants. The ego-vehicle’s plan is to drive from light to left while avoiding collisions with the other two obstacle vehicles. The prediction of the most probable mode of the obstacle vehicles is represented by colored ellipsoids. The driven trajectory by the ego-vehicle is depicted in green. The short-term planned trajectory is illustrated in purple.
@@
In this scenario, the ego vehicle is approaching a T-junction, with no traffic rules, in which its mission is to follow its designated lane while being uncertain about the intentions of the other vehicles as shown in Fig. 8. Here, we consider the case of a multi-vehicle traffic scenario in which two obstacle vehicles approach the T-junction where $\Lambda_{1}=\{\texttt{lane keep},\texttt{left turn}\}$, and $\Lambda_{2}=\{\texttt{yield},\texttt{left turn}\}$. In this case, it is not sufficient to consider the belief of a single agent’s intention as we did in the previous scenario, however, instead, we need to get a belief about how the traffic scene will evolve by considering all permutations the traffic participants can have as stated in (7).
@@ VI-C1 Environment Setup
Figure 9: The evolution of the belief of both obstacles based on their observed orientations according to (6).
In this scenario, a horizon of $N=25$ steps is defined, with a discretization step of 0.2 s, resulting in a time horizon of 5.0 s, and a branching time $t_{b}=2.4$ s is defined. As a benchmark comparison, we evaluated the same task using a robust planner that optimizes a single plan that considers all the modes that the other agents could have, and a greedy baseline that only considers the most probable predicted mode of each obstacle. All optimization parameters for both methods are set to be identical to guarantee a fair comparison. At the start of the simulation, all permutations, $\theta_{i}\in\Theta$, are initialized with equal likelihoods, $\theta_{i}=0.25$. The evolution of the belief over both obstacle modes is illustrated in Fig. 9.
@@
VI-C2 Quantitative Results
@@
Here it is worth pointing out that the proposed algorithm is implemented in Python to interface with CommonRoad. The average computational time over the experiments is 186 ms. However, since the sampling-based approach is parallelizable, this computational time can be further improved by evaluating the constraints of the sampled trajectories through parallelizable computations.
@@
VI-C3 Effect of branching time on the plan
In this section, an analysis of how the branching time affects contingency planning is conducted. For this purpose, we run experiments for all values of branching time $t_{b}\in[\Delta t,T]$, where $\Delta t$ is the discretization step that we set to 0.2 s. For each branching time, we run 100 simulated experiments in which the obstacles’ intents and their initial states are randomly initialized. We analyze the effect of the branching time on the relative average velocity the ego-vehicle exhibits with respect to the baseline, $t_{b}=T$. Moreover, the maximum risk among all experiments for each branching time is recorded. The results are reported in Fig. 10. As shown, for larger values of the branching time, the performance gap between both methods is small since most of the plan is constituted by the shared plan and thus the effect of the belief updater in the cost function is not pronounced. Indeed, when $t_{b}=T$, the disparity in performance disappears, as our proposed approach aligns with the baseline method under such conditions. For earlier branching times, however, the performance gap becomes more pronounced since the future information gain beyond the branching time is well exploited in the cost function because of the additional degrees of freedom introduced by the contingent plans. By inspecting the maximum risk plot depicted in Fig. 10, it can be observed that the maximum risk, $\eta$, increases as the branching time becomes shorter. This can be attributed to the over-confidence in the planned trajectory after the branching time causing the ego-vehicle to take more risky maneuvers. For sufficiently short branching times, $t_{b}\leq 2.0$ s in this example, the ego-vehicle could not find a feasible trajectory that does not violate the maximum risk, $\delta=0.05$, in the chance constraint, and subsequently, we execute the planned with the least risk that is dynamically feasible as we indicated earlier in Remark 3. This concludes that the branching time in contingency planning is related to the maximum risk the ego-vehicle perceives. More analysis regarding the estimation of the branching time $t_{b}$ based on the ego-vehicle’s dynamics capabilities is left for future work.
VI-C4 Branching time estimation

### v2

@@ II-A1 Planning with multi-modal predictions
Another line of work for planning under uncertainty is by formulating the planning problem as a partially observable Markov decision process (POMDP) by constructing a belief tree based on a discrete set of obstacle vehicle’s intentions [28, 29]. Solving such problems, however, becomes computationally intractable when the problem size scales. To address this problem, multi-policy decision-making [33, 34, 35] decomposes the belief tree into a limited number of closed-loop policies by leveraging semantic information. However, these approaches solve for the best ego-policy over all possible future realizations resulting in overly-conservative plans that do not exploit the multi-modality in the obstacle-vehicle behavior. To tackle this problem, inspired by branch-MPC, TPP [36] proposes an approach that converts the continuous space motion planning problem into a tractable problem by converting both the trajectory tree and scenario tree into a finite-horizon MDP. The optimal policy is then determined via dynamic programming over the constructed MDP. Despite its scalability to multiple vehicles, TPP outputs a single optimal policy over all scenarios causing unavoidable loss of multi-modality information.
This issue can be tackled by separating the planning problem into short-term and long-term responses [14, 15, 17, 16]. This is realized by generating a short-term trajectory, for $t<t_{b}$, where $t_{b}$ indicates a branching time, which guarantees safety with respect to all possible future intents that the other agents can exhibit. After $t_{b}$ duration, it is assumed that the ego vehicle will be able to determine the intent of the human driver, and thus it is sufficient to safeguard against the most likely intent. Thus, the contingent plan can ensure safety only when the ego vehicle acquires a clear understanding of the intentions of other agents by the time of branching. However, in these works, the probability associated with the uncertain intents of the dynamic agents is assumed to be fixed, and the planning is executed in an open-loop fashion. Thus, an exact estimation of the branching time $t_{b}$ is required, otherwise a collision may occur [17]. Learning-based approaches have been recently introduced in the contingency planning context [37, 38], however, they are not interpretable and hard to tune. [66] proposes a contingency planning approach that uses a dynamic branching point determined by a predefined heuristic. This heuristic chooses the branching time as the maximum time such that any two future scenarios starting at the current time only diverge by a maximum distance. Despite being effective, this heuristic entails at least double the computational time since the divergence measure is invoked on the ego-vehicle trajectories.
@@ V-A3 Probabilistic Risk Assessment
Here, we approximate the ego-vehicle by two discs, where each disc is referred to by $d$, and calculate the probability of collision for each disc. Note that we model the ego-vehicle as two discs such that we can consider all its shape in calculating the induced risk. Given the probability density function of the prediction model as indicated in (2), the probability of collision, for each ego-vehicle’s disc ate each planning stage $k$ per obstacle $o$, can be calculated by estimating the spatio-temporal overlap of the predicted modes with the ego-vehicle’s plan $\tau$,
@@
which is an integral of the Mixture of Gaussians (MoG) probability density function over a specified domain where the integration domain $D$ is defined as a circle whose center is located at the predicted vehicle pose $\boldsymbol{x}_{k}^{d}$ at stage $k$ along the prediction horizon, and its radius $r$ is the sum of the vehicle and obstacle radii since the PDF does not account for the obstacle-vehicle’s shape.
@@
After calculating the risk for each ego-vehicle’s disc at every timestep $k$ along the planning horizon per each obstacle $o$, the predicted risk $\eta$ at the current time step is defined by maximizing risk for all obstacles, $o\in\mathcal{I}_{o}$, at all planning horizon stages $k$ and all discs $d$,
@@
Figure 3: Snapshots from the CommonRoad simulation framework for an overtaking scenario where the ego-vehicle is depicted by the car icon and the obstacle-vehicle is illustrated as a blue rectangle. The obstacle-vehicle has three policies that it can execute, Λ={maintain speed,slow down,lane $\Lambda=\{\texttt{maintain speed},\texttt{slow down},\texttt{lane change}\}$, where the obstacle-vehicle’s predicted trajectories are illustrated in distinct colors in 3(a). We first sample a set of short-term plans, $\mathcal{T}_{0:t_{b}}$ visualized in black, till the branching time $t_{b}$ shown in 3(a). We then sample a set of long-term plans, $\mathcal{T}_{t_{b}:T}$ depicted in purple, conditioned on the terminal states of the short-term plans, shown in 3(b). For the long-term plans, a cost is assigned to each plan per obstacle-vehicle policy $\lambda_{i}$. In Fig. 3(c), the total cost of the entire plan is computed by (13a), and the total optimal plan is the one given by $\tau^{*}=\tau^{*}_{0:t_{b}}\cup\{\tau^{*}_{t_{b}:T}(\lambda_{1}),\tau^{*}_{t_{b}:T}(\lambda_{2}),\tau^{*}_{t_{b}:T}(\lambda_{3})\}$. In this example, the belief over the first policy is dominant, $b(\lambda_{1})=0.63$, compared to the rest causing the optimal short-term plan $\tau^{*}_{0:t_{b}}$ to be more biased towards $\tau^{*}_{t_{b:T}}(\lambda_{1})$.
@@ V-B Planner Design
Here it is crucial to emphasize that an advantage of using a sampling-based planner in the Frenet frame as a basis for our contingency planning framework is that the sampled trajectories cover various maneuvers that the ego-vehicle can have. Thus, in contrast to the approaches proposed in [30, 31, 32], our approach eliminates the necessity for a pre-established trajectory tree or branching topology to formulate contingent plans. Subsequently, the ego-vehicle is not confined to a specific set of predefined policies. This concept is further explained in Fig. 3 where we show how the optimal plan is selected from the sampled short-term and long-term trajectories. Additionally, it illustrates how the short-term plan is shaped based on the belief the ego-vehicle maintains over the obstacle-vehicle policies. Fig. 5 shows how the belief the ego-vehicle maintains over the long-term plans affects the behavior of the short-term plan. It is observed that when the ego-vehicle has a higher belief in one of the human’s intents, the shared plan tends to be biased towards the corresponding contingent plan. In particular, as illustrated in Fig. 5-left, when the ego-agent has a higher belief that the human-driven vehicle aims at executing the lane-keeping policy, the shared-plan biases its motion to accelerate along its lane. On the other hand, when the higher belief is assigned to the lane-change maneuver, the shared-plan tends towards steering a bit to the right while decelerating. This is achieved by balancing the cost of the shared-plan according to the likelihood of the beliefs.
@@ VI-A Experimental Setup
In contrast to the branch-MPC approach [30] which encounters scalability challenges when addressing multiple obstacle vehicles, primarily due to its exponential complexity requiring a pruning protocol, our proposed approach exhibits seamless adaptability to multi-vehicle scenarios. This is illustrated in the T-junction, and intersection scenarios, where the ego-vehicle interacts with multiple vehicles whose intents are not known to the ego-vehicle a priori [^1] A video of the simulated experiments accompanies this paper.. The computer running the simulations is equipped with an Intel® CoreTM i7 CPU@2.6GHz.
@@ VI-B1 Obstacle-vehicle Model
Non-cont. MLE [50] 3.71 (0.324) 17.62 (0.51) 0.48 (0.12) 2.74 % 0.1081 5.278 (6.13)
Contingent w/o belief updater [15] 4.46 (0.308) 16.68 (0.78) 0.84 (0.08) 0.00 % 0.0462 2.693 (4.63)
@@ VI-B3 Qualitative Results
Fig. 8 shows the evolution of the ego-vehicle’s belief over the obstacle-vehicle policies over time. At the beginning, the ego-vehicle reveals its lane-change intention by swerving into the obstacle’s vehicle lane. Due to the reactive behavior of the obstacle-vehicle, the probability of it executing a lane-change in the ego-vehicle’s lane drops. However, the ego-vehicle could not complete the lane change since no valid trajectory is obtained that does not violate the safety constraints, and thus returns to its original lane. Since the lane-change policy of the obstacle-vehicle, $b_{3}$, becomes relatively low, the ego-vehicle then initiates another attempt to overtake the obstacle vehicle which probes it to decelerate allowing the ego-vehicle to complete the lane change. To visualize how the obstacle-vehicle’s policy changes with the timesteps, Fig. 7 shows the evolution of the obstacle-vehicle’s velocity with the timesteps. As depicted, the ego-vehicle exhibits maintain speed policy till 2.7s. It then switches to slow down policy from 2.7s to 3.2s. This corresponds to the moment at which the ego-vehicle initiates its lane-change maneuver. It then switches back to the maintain speed policy.
Figure 7: The evolution of the obstacle-vehicle’s velocity with time
On the other hand, it was observed that the branch-MPC [30] brakes strangely while executing the lane-change maneuver relying on the obstacle-vehicle to yield to the ego-vehicle  [^2] The reader can refer to the video supplement, at 02:01, to observe such behavior.. However, such a maneuver is risky since it may result in collisions if the obstacle-vehicle does not react on time.
Figure 8: Belief evolution over obstacle vehicle’s policies along the timesteps. $b_{1},b_{2},b_{3}$ represent maintain fixed speed, slow down, and lane-change policies respectively. The gray vertical lines denote the time instants at which the snapshots in Fig. 6 are taken.
@@ VI-B4 Quantitative Results
Figure 9: Snapshots from the T-junction simulated environment in CommonRoad at different time instants. The ego-vehicle’s plan is to drive from right to left while avoiding collisions with the other two obstacle vehicles. In Fig 9(a), the prediction of both modes of both obstacles is represented by colored ellipsoids where the blue ellipsoids depict the left turn and yield policies for the leftward and upward vehicles respectively whereas the yellow ellipsoids illustrate the lane keep and left turn policies. The driven trajectory by the ego-vehicle is depicted in green. The short-term planned trajectory is illustrated in purple.
@@
In this scenario, the ego vehicle is approaching a T-junction, with no traffic rules, in which its mission is to follow its designated lane while being uncertain about the intentions of the other vehicles as shown in Fig. 9. Here, we consider the case of a multi-vehicle traffic scenario in which two obstacle vehicles approach the T-junction where $\Lambda_{1}=\{\texttt{lane keep},\texttt{left turn}\}$, and $\Lambda_{2}=\{\texttt{left turn},\texttt{yield}\}$. In this case, it is not sufficient to consider the belief of a single agent’s intention as we did in the previous scenario, however, instead, we need to get a belief about how the traffic scene will evolve by considering all permutations the traffic participants can have as stated in (7).
@@ VI-C1 Environment Setup
Figure 10: The evolution of the belief of both obstacles based on their observed orientations according to (6). Top: the belief evolution for obstacle 1 where $\lambda_{1}$ represents the lane keep policy and $\lambda_{2}$ represents the left turn policy. Bottom: the belief evolution for obstacle 2 where $\lambda_{1}$ represents the yield policy whereas $\lambda_{2}$ represents the left turn policy. Note that this represents one of the four permutations that the traffic scene can evolve to.
Figure 11: The evolution of the belief over permutations in the T-junction scenario where $\theta_{1}$ corresponds to lane keep and left turn for the leftward and upward vehicles respectively. The gray vertical lines denote the time instants at which the snapshots in Fig. 9 are taken.
In this scenario, a horizon of $N=25$ steps is defined, with a discretization step of 0.2 s, resulting in a time horizon of 5.0 s, and a branching time $t_{b}=2.4$ s is defined. As a benchmark comparison, we evaluated the same task using a robust planner that optimizes a single plan that considers all the modes that the other agents could have, and a greedy baseline that only considers the most probable predicted mode of each obstacle. All optimization parameters for both methods are set to be identical to guarantee a fair comparison. At the start of the simulation, all permutations, $\theta_{i}\in\Theta$, are initialized with equal likelihoods, $\theta_{i}=0.25$. The evolution of the belief over both obstacle modes is illustrated in Fig. 10. In Fig. 10, we present the dynamic evolution of the ego-vehicle’s belief regarding various intentions of surrounding obstacles over time. Specifically, Fig. 10(a) illustrates this evolution for a leftward-bound vehicle encountering a T-junction, where the obstacle vehicle faces the choice between continuing straight or executing a left turn. Similarly, Fig. 10(b) portrays the belief dynamics for an upward-bound vehicle confronted with the options of turning left or yielding to the ego-vehicle. In Fig. 10(a), we observe the ego-vehicle’s initial struggle with uncertainty regarding the leftward vehicle’s intentions, reflected in an equal belief distribution ($b(\lambda_{1})=b(\lambda_{2})=0.5$) as its state aligns with the mean of both distributions. However, as the obstacle vehicle’s state gradually deviates from this equilibrium, the belief over the left-turn maneuver diminishes, leading to a corresponding increase in belief regarding the alternative mode ($b(\lambda_{1})$). This nuanced adjustment allows the ego-vehicle to attenuate its emphasis on the left-turn possibility, thereby facilitating an accelerated trajectory within the T-junction. Analogous dynamics are observed for the second obstacle vehicle, as depicted in Fig. 10(b). Subsequently, the updated beliefs per mode, $\lambda\in\Lambda$, are utilized to update the probabilities over the different permutations, $\theta\in\Theta$, the traffic scene can evolve to as shown in Fig. 11. As illustrated in Fig. 11, the evolution of permutations depicts a gradual decrease in the belief regarding $\theta_{2}$ and $\theta_{3}$ over successive iterations, ultimately diminishing after approximately 58 time steps where their influence on the ego-vehicle’s plan is disregarded. Furthermore, by the 76th time-step, the ego-vehicle attains a high level of certainty that $\theta_{1}$ is the accurate hypothesis adopted by the obstacles. As a result, only this prediction mode significantly impacts the planner’s decision-making process.
@@
VI-C2 Prediction model
In this scenario, we use a synthesized prediction model that incorporates the multi-modality in the obstacle-vehicle’s intentions. Specifically, our multi-modal prediction model works as follows:
  1. (i)
    By identifying target lanes in the T-junction, we extrapolate the intended trajectories of surrounding vehicles.
  2. (ii)
    By employing a motion planner for each vehicle, we generate ground truth trajectories towards these lanes, resulting in multi-modal trajectories per vehicle.
  3. (iii)
    Our multi-modal prediction strategy involves:
    • Utilizing a uni-modal prediction model trained on extensive CommonRoad datasets [67] to generate Gaussian trajectory distributions for each potential mode, corresponding to the trajectories from Step (ii).
    • Amalgamating these distributions into a Gaussian Mixture Model (GMM), weighted by their likelihoods, inspired by prior works such as [19, 18].
  4. (iv)
    Mode weights, determining the likelihood of each trajectory mode, are computed based on collision avoidance metrics. These metrics, quantifying collision probabilities between obstacle vehicle trajectories and the ego-vehicle’s planned trajectory, dynamically adjust mode weights using a softmax function inspired by works such as [30].
Remark 6
We emphasize that our approach is agnostic to the prediction model employed. Any prediction model capable of providing Gaussian distributions over the predicted modes can be utilized.
VI-C3 Quantitative Results
@@
Here it is worth pointing out that the proposed algorithm is implemented in Python to interface with CommonRoad. The average computational time over the experiments is 174.98 ms. Tab. III shows how the computational time $t_{c}$ scales with the number of agents. We emphasize that since the sampling-based approach is parallelizable, the computational time can be further improved by evaluating the constraints of the sampled trajectories through parallelizable computations.
Number of agents 1 2 3
$t_{c}$ (ms) 134.53 186.12 204.31
$\sigma_{t_{c}}$ (ms) 4.7 6.9 11.8
TABLE III: Average computational time $t_{c}$ and standard deviation $\sigma_{t_{c}}$ for increasing number of agents.
@@
VI-C4 Effect of branching time on the plan
In this section, an analysis of how the branching time affects contingency planning is conducted. For this purpose, we run experiments for all values of branching time $t_{b}\in[\Delta t,T]$, where $\Delta t$ is the discretization step that we set to 0.2 s. For each branching time, we run 100 simulated experiments in which the obstacles’ intents and their initial states are randomly initialized. We analyze the effect of the branching time on the relative average velocity the ego-vehicle exhibits with respect to the baseline, $t_{b}=T$. Moreover, the maximum risk among all experiments for each branching time is recorded. The results are reported in Fig. 12. As shown, for larger values of the branching time, the performance gap between both methods is small since most of the plan is constituted by the shared plan and thus the effect of the belief updater in the cost function is not pronounced. Indeed, when $t_{b}=T$, the disparity in performance disappears, as our proposed approach aligns with the baseline method under such conditions. For earlier branching times, however, the performance gap becomes more pronounced since the future information gain beyond the branching time is well exploited in the cost function because of the additional degrees of freedom introduced by the contingent plans. By inspecting the maximum risk plot depicted in Fig. 12, it can be observed that the maximum risk, $\eta$, increases as the branching time becomes shorter. This can be attributed to the over-confidence in the planned trajectory after the branching time causing the ego-vehicle to take more risky maneuvers. For sufficiently short branching times, $t_{b}\leq 2.0$ s in this example, the ego-vehicle could not find a feasible trajectory that does not violate the maximum risk, $\delta=0.05$, in the chance constraint, and subsequently, we execute the planned with the least risk that is dynamically feasible as we indicated earlier in Remark 3.
(a) At $t=2.8$ s
(b) At $t=3.6$ s
(c) At $t=5.2$ s
Figure 13: Snapshots from the intersection simulated environment in CommonRoad at different time instants. The ego-vehicle’s plan is to take a left turn while avoiding collisions with the other two obstacle vehicles. The ego-vehicle is depicted as a car icon whereas the obstacle vehicles are represented by blue rectangles. The prediction of the most probable mode of the obstacle vehicles is represented by colored ellipsoids. The driven trajectory by the ego-vehicle is depicted in green. The short-term planned trajectory is illustrated in purple.
This concludes that the branching time in contingency planning is related to the maximum risk the ego-vehicle perceives. More analysis regarding the estimation of the branching time $t_{b}$ based on the ego-vehicle’s dynamics capabilities is left for future work.
VI-C5 Branching time estimation